Can you help? x=-5 what is the slope. Thank you! Shelly
Found 2 solutions by MathLover1, stanbon:
Answer by MathLover1(20855)

x=-5
there is no slope, we have a vertical line at x=-5


Answer by stanbon(75887)

Can you help? x=-5 what is the slope. Thank you! Shelly
--------------
If x = -5, 2 of the points might be (-5,0) and (-5,1)
---
slope = (1-0)/(-5--5) = 1/0 which is undefiined.
---------------
The line is a vertical line (straight up and down).
The slope of any vertical line is undefined.
----
Cheers,
Stan H.